OpenStudy (superhelp101):

Which of the following is produced when propane (C3H8) is combusted completely? (2 points) H2O CO O2 H2

OpenStudy (photon336):

Note complete combustion excess of oxygen, is different from incomplete combustion which would produce carbon monoxide. The formula for complete combustion is this: C3H8 + O2 = CO2 + H2O H2O would be your answer

OpenStudy (photon336):

CO is the product of incomplete combustion. While O2 is your reactant, can't be your product just doesn't make sense.

OpenStudy (photon336):

Apologies that formula above isn't balanced It should be this C3H8 + 5O2 = 3CO2 + 4H2O